How do you determine rather to add or to subtract when you move the variable by undoing it? Example:

11a - 3=5a + 9

Answer by nyc_function(2741)

11a - 3=5a + 9


When crossing the equal symbol, the signs of the terms and constants change to the opposite value.


We solve the equation by combining like terms.


11a - 5a = 9 + 3


Do you see that 5a became -5a and -3 became 3?


6a = 12

a = 12/6

a = 2